# Tiling spaces are covering spaces over irrational tori

### v1

@@ 3.2 The bundle $\hat{p}:\Omega\to\mathbb{T}_{\alpha}$
The first approximant [^2] We omit the technicality of whether the substitution rule forces the border because both complexes $\Omega_{0}$ and $\Omega_{1}$ from Anderson-Putnam construction [2] work for our purpose. $\Omega_{0}$ is the wedge of two circles with lengths 1 and $\alpha$, one for each letter of the alphabet $\{0,1\}$, the following approximants $\Omega_{n}$ are the same wedge of circles whose radii are $\alpha$ times the radii of the previous and subdivided by the substitution rule, where $\alpha$ is the expansion factor of the substitution rule $\omega$ (see Theorem 2 and Fig. 1). The transition functions $f_{n}$ forget the tiles from the $n$-th level. The general case for substitution rules with more than two letters is treated in the same fashion.
The space $\Omega$ is a (topological) fiber bundle over the one-dimensional torus $\mathbb{S}^{1}$ and its fibers are homeomorphic to the Cantor set $K$ (see Theorem 3). Consider such fibration
@@
In the diffeological context, it is perhaps more natural to consider a similar map $\hat{p}:\Omega\to\mathbb{T}_{\alpha}$ over the irrational torus $\mathbb{T}_{\alpha}=\mathbb{R}/(\mathbb{Z}+\alpha\mathbb{Z})$, which may be defined by mapping the vertices of a tiling $T$ to $\mathbb{Z}+\alpha\mathbb{Z}$ instead of $\mathbb{Z}$. We define $\hat{p}$ as the composition of maps
$\Omega\stackrel{{\scriptstyle\pi_{0}}}{{\longrightarrow}}\Omega_{0}\stackrel{{\scriptstyle\hat{q}}}{{\longrightarrow}}\mathbb{R}\stackrel{{\scriptstyle\mathrm{pr}_{\alpha}}}{{\longrightarrow}}\mathbb{T}_{\alpha},$
in this case $\pi_{0}:\Omega\to\Omega_{0}$ is the same projection as before, $\hat{q}$ is the function that maps the circles forming $\Omega_{0}$ onto the intervals of the form $[n+m\alpha,n+1+m\alpha]$ or $[n+m\alpha,n+(m+1)\alpha]$ for $n,m\in\mathbb{Z}$, and $\mathrm{pr}_{\alpha}$ is the quotient map $\mathrm{pr}_{\alpha}:\mathbb{R}\to\mathbb{R}/(\mathbb{Z}+\alpha\mathbb{Z})$. Thus $\hat{p}$ is the function that maps a tiling $T$ to the location of its origin into intervals like $[-1,0],[0,\alpha]\subset\mathbb{R}$, followed by the quotient map $\mathrm{pr}_{\alpha}$.
We endow $\Omega_{0}$ with the push-forward diffeology ${\pi_{0}}_{*}(\mathcal{D})$ of $\mathcal{D}$ through the map $\pi_{0}$, $\mathbb{R}$ with the standard diffeology and $\mathbb{T}_{\alpha}$ with the quotient diffeology for $\mathbb{R}/(\mathbb{Z}+\alpha\mathbb{Z})$.
@@
With the aforementioned diffeologies the map $\hat{p}:\Omega\to\mathbb{T}_{\alpha}$ is a smooth surjection.
@@
The map $\hat{p}$ is clearly surjective as, for a given $z\in\mathbb{T}_{\alpha}$ represented by $t+n+\alpha m$ with $t\in\mathbb{R},n,m\in\mathbb{Z}$, there is a tiling $T\in\Omega$ whose origin lies in $t\mod 1$ that is mapped to $z$. For proving smoothness note that the maps $\pi_{0}$ and $\mathrm{pr}_{\alpha}$ are smooth by definition and $\hat{q}\circ\pi_{0}$ is locally smooth in every tiling $T$ whose origin is not located in an end-point of a tile, because there are $D$-open intervals $A\subset\Omega$ where the plots parametrize translations within $A$. That is, given a plot $P$ of $\Omega$, $\hat{q}\circ\pi_{0}\circ P$ is a translation in $\mathbb{R}$, which is smooth and thus a plot. Hence $\hat{p}=\mathrm{pr}_{\alpha}\circ\hat{q}\circ\pi_{0}$ is locally smooth in every tiling $T$ whose origin is not located in the border of a tile.
However $\hat{p}$ is also locally smooth also in the rest of the tilings because the endpoints of the tiles are smoothly identified in $\mathbb{T}_{\alpha}$, by definition of its diffeology. Thus $\hat{p}$ is locally smooth in every point of $\Omega$ so it is smooth.
3.2.1 The structure groupoid for $\hat{p}:\Omega\to\mathbb{T}_{\alpha}$.
We proceed to investigate the structure groupoid associated to the smooth surjection $\hat{p}:\Omega\to\mathbb{T}_{\alpha}$. Recall from 2.2.2. that $G_{0}=\mathbb{T}_{\alpha}$ and
@@
for $\Omega_{z}:=\hat{p}^{-1}(z)$. The maps source $s:G_{1}\to G_{0}$ and target $t:G_{1}\to G_{0}$ are defined by $s(\varphi)=\hat{p}(dom(\varphi))$ and $t(\varphi)=\hat{p}(ran(\varphi))$.
@@
The map $\hat{p}:\Omega\to\mathbb{T}_{\alpha}$ is a diffeological fibration.
@@
We must show that the structure groupoid $(G_{1},G_{0})$ associated to $\hat{p}$ is fibrant, i.e. the characteristic map $\chi:G_{1}\to\mathbb{T}_{\alpha}\times\mathbb{T}_{\alpha}$ is a subduction [^3] See discussion after Example 2.. In order to conclude that, we have to show that for every plot $Q:U\to\mathbb{T}_{\alpha}\times\mathbb{T}_{\alpha}$ and every point $u\in U$ there exists a neighborhood $V$ of $u$ such that either $Q|_{V}$ is constant or there exists a plot $P:V\to G_{1}$ that locally lifts $Q$.
@@
where $\Omega_{s\circ P}$ stands for the pull-back of $\hat{p}$ along $s\circ P$, and analogously $\Omega_{t\circ P}$, are smooth.
Let $Q:U\to\mathbb{T}_{\alpha}\times\mathbb{T}_{\alpha}$ be a plot for the product diffeology on $\mathbb{T}_{\alpha}\times\mathbb{T}_{\alpha}$, then $Q(u)=(Q_{1}(u),Q_{2}(u))$ where $Q_{i}$ a plot of $\mathbb{T}_{\alpha}$. Moreover each $Q_{i}$ lifts locally (say in neighborhoods $V_{i}$, with $V_{1}\times V_{2}:=V\subset U$) to a plot $\bar{Q}_{i}$ of $\mathbb{R}$ because the diffeology on $\mathbb{T}_{\alpha}$ coincides with the quotient diffeology $\mathbb{R}/(\mathbb{Z}+\alpha\mathbb{Z}$).
Let $u\in U$, $z_{s}=\bar{Q}_{1}(u)$, $z_{t}=\bar{Q}_{2}(u),\bar{z}_{s}=\mathrm{pr}_{\alpha}(z_{s})$ and $\bar{z}_{t}=\mathrm{pr}_{\alpha}(z_{t})$. Then $z_{s},z_{t}\in\mathbb{R}$ and $\bar{z}_{s},\bar{z}_{t}\in\mathbb{T}_{\alpha}$. Let $\varphi_{z_{s},z_{t}}:\Omega\to\Omega$ be the map given by
@@
Note that this map translates the fiber $\hat{p}^{-1}(\bar{z}_{s}):=\Omega_{\bar{z}_{s}}$ to $\hat{p}^{-1}(\bar{z}_{t}):=\Omega_{\bar{z}_{t}}$. It is moreover a diffeomerphism between these fibers: it is smooth because it is a translation as well as its inverse. Then it is an element of $G_{1}$ satisfying $s(\varphi_{z_{s},z_{t}})=\bar{z}_{s}$ and $t(\varphi_{z_{s},z_{t}})=\bar{z}_{t}$. Thus we define a parametrization $P:V\to G_{1}$ given by $P(u)=\varphi_{z_{s},z_{t}}$ and assert that it is the plot we needed.
The maps $P_{s}$, $P_{t}$ are immediately seen to be smooth because they are translations of tilings, and $\chi\circ P$ equals $(\mathrm{pr}_{\alpha}\times\mathrm{pr}_{\alpha})\circ\bar{Q}$ which is a composition of smooth maps, and so it is also smooth. Hence $P$ is a plot of $G_{1}$ and it lifts the plot $Q:U\to\mathbb{T}_{\alpha}\times\mathbb{T}_{\alpha}$ by construction. Consequently $\chi$ is a subduction and the groupoid $(G_{1},G_{0})$ is fibrant.
@@
Let $\Omega$ be an unidimensional tiling space of primitive substitution tilings with a rule $\omega$ and expansion factor $\alpha$. Then $\Omega$ is a covering space over the irrational torus $\mathbb{T}_{\alpha}$.
@@
This map has the advantage over Williams’ and Sadun-Williams constructions that the proportion of the length of the tiles is preserved. So it provides a necessary condition for diffeomorphism between tiling spaces and thus for orbital equivalence. This due to the classification of irrational tori [17], as we shall see in the following section.
Remark 3.
This result is fairly general within the class of one-dimensional substitutional tiling spaces. In this case, there is not too much room for distinction of tilings (as in their higher dimensional counterparts) other than the combinatorics and lenghts of the tiles. This information is pretty much encoded in the dynamics of the substitution matrix. There is a close relation between these dynamics with the arithmetics of the eigenvalues of this matrix and the possibility of obtaining the same tilings by means of a cut-and-project scheme.
In fact, one-dimensional tilings lie within the intersection of the classes of substitution and projection tilings if the irrational number $\alpha$ is a quadratic irrational Pisot number [^4] A real number $\alpha$ is a Pisot-Vijayaraghavan number, also called PV-number or just Pisot number, if it is the root of a monic polynomial with integer coeficients, $\alpha>1$, and all of its Galois conjugates are less than 1 in absolute value. [16, 27] and the points in the border of the window $W$ are elements of $\mathbb{Q}[\alpha]$. In the next section we will turn our attention to the tiling spaces arising from the cut-and-project method.
@@ 4 Diffeological and sturmian equivalences for Tiling spaces
In the present section we study another bundle structure from projection method tiling spaces $\Omega_{\alpha}$ over $\mathbb{T}_{\alpha}$. We will see how to bring notions of equivalences between sturmian systems (Lemma 1) into the context of tiling spaces through their diffeological structures.
As discussed in Section 2.1.4 above, we can relate explicitly the sturmian sequences of $(X_{\alpha},\sigma)$ with canonical projection tilings in $\Omega_{\alpha}$. We define a function $\Phi:\Omega_{\alpha}\to X_{\alpha}$ which“forgets” the length of the tiles but preserves the combinatorics of every tiling, i.e. the funcion $\Phi$ maps the tiling $T$ onto the sequence $S\in\{0,1\}^{\mathbb{Z}}$ encoding the tiles of $T$, where $\{0,1\}$ represent the proto-tile set of $T$.
This function is not invertible but there is a canonical section $\Psi:X_{\alpha}\to\Omega_{\alpha}$ which assigns to a sturmian sequence $S$ the tiling $\Psi(S)=T$ in the tile set $\{0,1\}$ with lenghts $\frac{1}{\sqrt{1+\alpha^{2}}}$ and $\frac{\alpha}{\sqrt{1+\alpha^{2}}}$, respectively, appearing with the same combinatorics as in $S$, and with its origin placed on the left vertex of the tile associated to $s_{0}$. We define a diffeology for $X_{\alpha}$ as the pull-back diffeology of $\Omega_{\alpha}$ along the map $\Psi$. Thus the map $\Psi$ is automatically smooth.
The following lemmas verify that we can bring notions of sturmian systems into the diffeological context.
@@
Let $P:U\to\Omega_{\alpha}$ a plot of $\Omega_{\alpha}$. We must show that $\Phi\circ P:U\to X_{\alpha}$ is a plot of $X_{\alpha}$. But this is the case if and only if $\Psi\circ\Phi\circ P$ is a plot of $\Omega_{\alpha}$. As $P$ is a plot of $\Omega_{\alpha}$, by definition of the generating family $\mathcal{F}$ (see Definition 12) for the diffeology of $\Omega_{\alpha}$, for every $u\in U$ there exists a neighborhood $V\subset U$ containing $u$, a smooth map $g:V\to W\subset\mathbb{R}$, and an element $F_{T}:W\to\Omega_{\alpha}$ of the generating family, with $T\in\Omega_{\alpha}$, such that $P|_{V}=F_{T}\circ g$. Then
$(\Psi\circ\Phi\circ P)|_{V}(v)=(\Psi\circ\Phi)\circ F_{T}\circ g(v)=(\Psi\circ\Phi)(T+g(v)).$ (15)
@@
We must show first that the composition of every plot of $X_{\alpha}$ with the shift map is a plot of $X_{\alpha}$ itself. Let $P:U\to X_{\alpha}$ be a plot of $X_{\alpha}$. By definition of the pull-back diffeology along the map $\Psi$ defined above, the composition $\Psi\circ P$ is a plot of $\Omega_{\alpha}$. Then, by definition of the generating family $\mathcal{F}$ (see Definition 12) for the diffeology of $\Omega_{\alpha}$, for every $u\in U$ there exists a neighborhood $V\subset U$ containing $u$, a smooth map $g:V\to W\subset\mathbb{R}$, and an element $F_{T}:W\to\Omega_{\alpha}$ of the generating family, with $T\in\Omega_{\alpha}$, such that
$(\Psi\circ P)|_{V}=F_{T}\circ g.$ (16)
@@
Thus,
$(\Psi\circ\sigma\circ P)|_{V}(v)=F_{T+\delta_{\alpha}}+g(v),$
and $\Psi\circ\sigma\circ P$ is a plot of $\Omega_{\alpha}$. Which means that $\sigma\circ P$ is a plot of $X_{\alpha}$. Therefore $\sigma$ is smooth.
@@
An inductive argument gives us the following
Corollary 1.
The $\mathbb{Z}$ action on $X_{\alpha}$ given by
@@
Let $X_{\alpha}$ and $X_{\beta}$ sturmian spaces endowed with the diffeology from above. They are orbit equivalent (Definition 7) if and only if they are diffeomorphic.
@@
The result is a direct consequence of Proposition 6, Lemmas 6 and 7, and the definition of the mappings $\Psi$ and $\Phi$. The actions may be seen in the following diagram
@@
where $\nu_{\alpha}:\Psi_{\alpha}\circ\sigma\circ\Phi_{\alpha}$, $\nu_{\beta}:\Psi_{\beta}\circ\sigma\circ\Phi_{\beta}$ and $g=\Psi_{\beta}\circ f\circ\Phi_{\alpha}$.
@@
Let $\Omega_{\alpha}$ a tiling space of canonical projection of dimension 2 to 1 and irrational slope $\alpha$. Then $\Omega_{\alpha}$ is a $\mathbb{R}$-principal bundle over the double pointed irrational tori $\mathbb{T}_{\alpha}^{\bullet}$.
@@
The function $\Psi$ is a bijection between the orbits of the shift $\sigma$ in $X_{\alpha}$ and the orbits by continuous translation in $\Omega_{\alpha}$. Thus we will prove that the the surjection $X_{\alpha}\to\mathbb{T}_{\alpha}$ given by $S(\rho)\mapsto\mathrm{pr}_{\alpha}(\rho)$, where $\mathrm{pr}_{\alpha}$ is defined by 7. Consider $\rho,\rho^{\prime}\in(0,1)$ such that
@@
and that $S(\rho),S(\rho^{\prime})$ represent the sequences $\{s_{n}\}_{n\in\mathbb{Z}},\{s^{\prime}_{n}\}_{n\in\mathbb{Z}}$ respectively. Thus, we have the following equalities

### v3

@@ 3.2 The bundle $\tilde{p}:\Omega\to\mathbb{T}_{\alpha}$
The construction of Theorem 3 relies on Theorem 2 as it uses the maps between $\Omega$ and its aproximant $\Omega_{0}$ in order to obtain a (topological) fiber bundle with total space $\Omega$, the one-dimensional torus $\mathbb{S}^{1}$ as base space, and fibers homeomorphic to the Cantor set $K$. Consider such fibration
@@
In the diffeological context, it is more natural to consider the following alternative map: let $\{a_{i}\}$ be the set of lengths of the tiles of $T$. Then, instead of the previous quotient map $\pi:\mathbb{R}\to\mathbb{S}^{1}$, we map onto the quotient space $\mathbb{R}/(a_{1}\mathbb{Z}+\ldots+a_{n}\mathbb{Z})$, by mapping the vertices of a tiling $T$ to $a_{1}\mathbb{Z}+\ldots+a_{n}\mathbb{Z}$, instead of $\mathbb{Z}$. Note that this is precisely a return module.
Definition 13.
Let $P$ be a patch of a tiling $T$ of $\mathbb{R}^{d}$. Define the return vectors of $P$ as the set of vectors $v$ mapping $P$ to other occurences of $P$ in the tiling, i.e. the vectors $v$ such that the patern at $P+v$ coincides with $P$. Define the return module $R_{P}$ of the patch $P$ as the $\mathbb{Z}$-span of the return vectors.
So, for a given tiling space $\Omega$ in $\mathbb{R}^{d}$, we pick a tiling $T\in\Omega$ and a vertex $v$ on it. Let $\mathbb{T}_{v}=\mathbb{R}^{d}/R_{v}$, where $R_{v}$ is the return module of the vertex $v$. We define $\tilde{p}$ as the composition of maps
$\Omega\stackrel{{\scriptstyle\hat{q}}}{{\longrightarrow}}\mathbb{R}^{d}\stackrel{{\scriptstyle\mathrm{pr}_{v}}}{{\longrightarrow}}\mathbb{T}_{v}$
In this case $\hat{q}$ is the function that maps a tiling $T$ onto the position of its origin in $\mathbb{R}^{d}$, and $\mathrm{pr}_{v}$ is the quotient map $\mathrm{pr}_{v}:\mathbb{R}^{d}\to\mathbb{T}_{v}$. Thus $\tilde{p}$ is the function that maps a tiling $T$ to the equivalence class of the location of its origin in the quotient space $\mathbb{T}_{v}$. Note that in the familiar case where $d=1$ and the tile set is $\{1,\alpha\}$, the return module $R_{v}$ coincides with $\mathbb{Z}+\alpha\mathbb{Z}$ and so the base space of the function $\tilde{p}$ is the irrational torus $\mathbb{T}_{\alpha}$.
@@
With the diffeology $\mathcal{D}$ for $\Omega$ from (12), the standard diffeology for $\mathbb{R}^{d}$, and the quotient diffeology for $\mathbb{T}_{v}$, the map $\tilde{p}:\Omega\to\mathbb{T}_{v}$ is a smooth surjection.
@@
The map $\tilde{p}$ is clearly surjective as, for a given $\bar{z}\in\mathbb{T}_{v}$ represented by $z+\sum_{i=1}^{n}v_{i}n_{i}$ with $z\in\mathbb{R},n_{i}\in\mathbb{Z}$, and $\{v_{i}\}$ the return vectors of a given vertex in $\Omega$, there is a tiling $T\in\Omega$ whose origin lies in $z$ that is mapped to $\bar{z}$. For proving smoothness note that the map $\mathrm{pr}_{v}$ is smooth by definition and every plot $P$ of $\Omega$ is locally a translation, so $\hat{q}\circ P$ is also a translation in $\mathbb{R}^{d}$ and translations are smooth, so it is a plot of $\mathbb{R}^{d}$. Hence $\tilde{p}$ is smooth. ∎
3.2.1 The structure groupoid for $\tilde{p}:\Omega\to\mathbb{T}_{v}$.
We proceed to investigate the structure groupoid associated to the smooth surjection $\tilde{p}:\Omega\to\mathbb{T}_{v}$. Recall from 2.2.2. that $G_{0}=\mathbb{T}_{v}$ and
@@
for $\Omega_{z}:=\tilde{p}^{-1}(z)$. The maps source $s:G_{1}\to G_{0}$ and target $t:G_{1}\to G_{0}$ are defined by $s(\varphi)=\tilde{p}(dom(\varphi))$ and $t(\varphi)=\tilde{p}(ran(\varphi))$.
@@
The map $\tilde{p}:\Omega\to\mathbb{T}_{v}$ is a diffeological fibration.
@@
We must show that the structure groupoid $(G_{1},G_{0})$ associated to $\tilde{p}$ is fibrating, i.e. the characteristic map $\chi:G_{1}\to\mathbb{T}_{v}\times\mathbb{T}_{v}$ is a subduction [^2] See discussion after Example 2.. In order to conclude that, we have to show that for every plot $Q:U\to\mathbb{T}_{v}\times\mathbb{T}_{v}$ and every point $u\in U$ there exists a neighborhood $V$ of $u$ such that either $Q|_{V}$ is constant or there exists a plot $P:V\to G_{1}$ that locally lifts $Q$.
@@
where $\Omega_{s\circ P}$ stands for the pull-back of $\tilde{p}$ along $s\circ P$, and analogously $\Omega_{t\circ P}$, are smooth.
Let $Q:U\to\mathbb{T}_{v}\times\mathbb{T}_{v}$ be a plot for the product diffeology on $\mathbb{T}_{v}\times\mathbb{T}_{v}$, then $Q(u)=(Q_{1}(u),Q_{2}(u))$ where $Q_{i}$ a plot of $\mathbb{T}_{v}$. Moreover each $Q_{i}$ lifts locally (say in neighborhoods $V_{i}$, with $V_{1}\times V_{2}:=V\subset U$) to a plot $\bar{Q}_{i}$ of $\mathbb{R}^{d}$ because the diffeology on $\mathbb{T}_{v}$ coincides with the quotient diffeology $\mathbb{R}^{d}/R_{v}$.
Let $u\in U$, $z_{s}=\bar{Q}_{1}(u)$, $z_{t}=\bar{Q}_{2}(u),\bar{z}_{s}=\mathrm{pr}_{v}(z_{s})$ and $\bar{z}_{t}=\mathrm{pr}_{v}(z_{t})$. Then $z_{s},z_{t}\in\mathbb{R}^{d}$ and $\bar{z}_{s},\bar{z}_{t}\in\mathbb{T}_{v}$. Let $\varphi_{z_{s},z_{t}}:\Omega\to\Omega$ be the map given by
@@
Note that this map translates the fiber $\tilde{p}^{-1}(\bar{z}_{s}):=\Omega_{\bar{z}_{s}}$ to $\tilde{p}^{-1}(\bar{z}_{t}):=\Omega_{\bar{z}_{t}}$. It is moreover a diffeomorphism between these fibers: it is smooth because it is a translation as well as its inverse. Then it is an element of $G_{1}$ satisfying $s(\varphi_{z_{s},z_{t}})=\bar{z}_{s}$ and $t(\varphi_{z_{s},z_{t}})=\bar{z}_{t}$. Thus we define a parametrization $P:V\to G_{1}$ given by $P(u)=\varphi_{z_{s},z_{t}}$ and assert that it is the plot we needed.
The maps $P_{s}$, $P_{t}$ are immediately seen to be smooth because they are translations of tilings, and $\chi\circ P$ equals $(\mathrm{pr}_{v}\times\mathrm{pr}_{v})\circ\bar{Q}$ which is a composition of smooth maps, and so it is also smooth. Hence $P$ is a plot of $G_{1}$ and it lifts the plot $Q:U\to\mathbb{T}_{v}\times\mathbb{T}_{v}$ by construction. Consequently $\chi$ is a subduction and the groupoid $(G_{1},G_{0})$ is fibranting.
@@
Let $\Omega$ be a space of repetitive tilings of $\mathbb{R}^{d}$. Then $\Omega$ is a covering space over $\mathbb{R}^{d}/R_{v}=\mathbb{T}_{v}$.
@@
This map has the advantage over Williams’ and Sadun-Williams constructions that the proportion of the length of the tiles is preserved. In the one-dimensional case, it provides a necessary condition for diffeomorphism between tiling spaces and thus for orbital equivalence. This is due to the classification of irrational tori [21], as we shall see in the following section.
@@ 4 Diffeological and sturmian equivalences for Tiling spaces
In the present section we study another bundle structure from projection method one-dimensional tiling spaces $\Omega_{\alpha}$ over $\mathbb{T}_{\alpha}$. We will bring notions of equivalence between sturmian systems (Lemma 1) into the context of tiling spaces through their diffeological structures. We will relate explicitly the sturmian sequences of $(X_{\alpha},\sigma)$ with canonical projection tilings in $\Omega_{\alpha}$.
For this matter we define a function $\Phi:\Omega_{\alpha}\to X_{\alpha}$ which “forgets” the length of the tiles but preserves the combinatorics of every tiling, i.e. the funcion $\Phi$ maps the tiling $T$ onto the sequence $S\in\{0,1\}^{\mathbb{Z}}$ encoding the tiles of $T$, where $\{0,1\}$ represent the proto-tile set of $T$. This function is not invertible but there is a canonical section $\Psi:X_{\alpha}\to\Omega_{\alpha}$ which assigns to a sturmian sequence $S$ the tiling $\Psi(S)=T$ in the tile set $\{0,1\}$ with lenghts $\frac{1}{\sqrt{1+\alpha^{2}}}$ and $\frac{\alpha}{\sqrt{1+\alpha^{2}}}$, respectively, appearing with the same combinatorics as in $S$, and with its origin placed on the left vertex of the tile associated to $s_{0}$. We define a diffeology for $X_{\alpha}$ as the pull-back diffeology of $\Omega_{\alpha}$ along the map $\Psi$. Thus the map $\Psi$ is automatically smooth.
@@
Let $P:U\to\Omega_{\alpha}$ be a plot of $\Omega_{\alpha}$. We must show that $\Phi\circ P:U\to X_{\alpha}$ is a plot of $X_{\alpha}$. But this is the case if and only if $\Psi\circ\Phi\circ P$ is a plot of $\Omega_{\alpha}$. As $P$ is a plot of $\Omega_{\alpha}$, by definition of the generating family $\mathcal{F}$ (see Definition 12) for the diffeology of $\Omega_{\alpha}$, for every $u\in U$ there exists a neighborhood $V\subset U$ containing $u$, a smooth map $g:V\to\mathbb{R}$, and an element $\hat{T}:\mathbb{R}^{d}\to\Omega_{\alpha}$ of the generating family, with $T\in\Omega_{\alpha}$, such that $P|_{V}=\hat{T}\circ g$. Then
$(\Psi\circ\Phi\circ P)|_{V}(v)=(\Psi\circ\Phi)\circ\hat{T}\circ g(v)=(\Psi\circ\Phi)(T+g(v)).$ (15)
@@
We must show first that the composition of every plot of $X_{\alpha}$ with the shift map is a plot of $X_{\alpha}$ itself. Let $P:U\to X_{\alpha}$ be a plot of $X_{\alpha}$. By definition of the pull-back diffeology along the map $\Psi$ defined above, the composition $\Psi\circ P$ is a plot of $\Omega_{\alpha}$. Then, by definition of the generating family $\mathcal{F}$ for the diffeology of $\Omega_{\alpha}$, for every $u\in U$ there exists a neighborhood $V\subset U$ containing $u$, a smooth map $g:V\to W\subset\mathbb{R}$, and an element $\hat{T}:\mathbb{R}^{d}\to\Omega_{\alpha}$ of $\mathcal{F}$, with $T\in\Omega_{\alpha}$, such that
$(\Psi\circ P)|_{V}=\hat{T}\circ g.$ (16)
@@
Then
$(\Psi\circ\sigma\circ P)|_{V}(v)=\widehat{(T+\delta_{\alpha})}(g(v)),$
thus $\Psi\circ\sigma\circ P$ is a plot of $\Omega_{\alpha}$. Which means that $\sigma\circ P$ is a plot of $X_{\alpha}$. Therefore $\sigma$ is smooth.
@@
An inductive argument gives us that the $\mathbb{Z}$-action on $X_{\alpha}$ given by
@@
Let $X_{\alpha}$ and $X_{\beta}$ be sturmian spaces endowed with the diffeology from above. They are orbit equivalent (Definition 7) if and only if they are diffeomorphic.
@@
The result is a direct consequence of Proposition 5, Lemmas 5 and 6, and the definition of the mappings $\Psi$ and $\Phi$. The actions may be seen in the following diagram
@@
where $\nu_{\alpha}:=\Psi_{\alpha}\circ\sigma\circ\Phi_{\alpha}$, $\nu_{\beta}:=\Psi_{\beta}\circ\sigma\circ\Phi_{\beta}$ and $g=\Psi_{\beta}\circ f\circ\Phi_{\alpha}$.
@@
Let $\Omega_{\alpha}$ a tiling space of canonical projection of dimension 2 to 1 and irrational slope $\alpha$. Then $\Omega_{\alpha}$ is a $\mathbb{R}$-principal bundle over the irrational tori with two origins $\mathbb{T}_{\alpha}^{\bullet}$ defined below.
@@
The function $\Psi$ is a bijection between the orbits under the shift $\sigma$ in $X_{\alpha}$ and the orbits by continuous translation in $\Omega_{\alpha}$. We prove first that the surjection $X_{\alpha}\to\mathbb{T}_{\alpha}$ given by $S(\rho)\mapsto\mathrm{pr}_{\alpha}(\rho)$, for $\mathrm{pr}_{\alpha}$ defined in 7, is well defined. Consider $\rho,\rho^{\prime}\in(0,1)$ such that
@@
and let $S(\rho),S(\rho^{\prime})$ represent the sequences $\{s_{n}\}_{n\in\mathbb{Z}},\{s^{\prime}_{n}\}_{n\in\mathbb{Z}}$ respectively. Thus, we have the following equalities